A cylindrical jar of diameter 14 cm and depth 20 cm is filled with water.when 300 lead shots of the same size are dropped into the jar.the level of water rises by 2.8 cm .find the diameter of each lead shot
Answer by mananth(16946)

A cylindrical jar of diameter 14 cm and depth 20 cm is filled with water.when 300 lead shots of the same size are dropped into the jar.the level of water rises by 2.8 cm .find the diameter of each lead shot
..
volume of the cylinder = pi*r^2*h . r=7
volume of water per cm
=pi*7^2*1
= 154 cu.cm
The increase in height = 2.8 cm
the volume increase= 154*2.8= 431.2 cu. cm.
..
The displaced water is the volume of the 1000 lead shots.
so volume of one shot = 431.2/1000
= 0.431 cu. cm.
..
Volume of lead shot = 0.431
volume = 4/3 *pi*r^3
0.431*3/4*pi= r^3
r^3= 0.1029
r= cube root (0.1029)
r=0.47 cm